100 g of manganese dioxide reacts with 300 g of hydrochloric acid in the reaction shown below.

MnO2 + 4HCl -> MnCl2+ 2H2O + Cl2 

 (Use the following molar masses: MnO2= 86.9g/mol HCl = 36.5g/mol ; MnCl2= 125.8g/mol; H2O= 18.02g/mol and Cl 2 =70.9 g/mol.

 a. Which reactant is limiting?

b. Which reactant is excess? 

c. What is the theoretical yield of Cl2? 

d. What is the percent yield of the reaction if 75.0 g of Cl2 was produced


1
Expert's answer
2022-04-06T14:20:03-0400

a.      Which reactant is limiting?

Moles MnO2:

100 / 86.9 = 1.15 (mol)

Moles HCl:

300 / 36.5 = 8.22 (mol)

The mole ratio of the reactants is 1 : 4. So, in order for 1.15 mol of MnO2 to reacts 1.15*4 = 4.6 mol of HCl required. Thus, MnO2 is limiting reagent.

 

b.      Which reactant is excess? 

HCl needed for the reaction is 4.6 mol. We’re given 8.22. So, HCl is in excess.

 

c.      What is the theoretical yield of Cl2

The ratio between MnO2 and Cl2 is 1 : 1. So, moles of Cl2 produced is 1.15 mol – theoretical yield.

Mass = 1.15 * 70.9 = 81.535 g.

 

d.      What is the percent yield of the reaction if 75.0 g of Cl2 was produced?

Percent yield = 75.0 / 81.535 = 0.92 or 92%.